1/1

N. N. 43,

An Irish Christmas Blessing

God grant you lightness in your step, a smile on every face you meet, Loved ones gathered at your hearth, and at your door, good friends to greet A holy hymn upon your lips, a window candle burning bright, And may the Good Lord bless your Heart and come to dwell there at Christmas night.

Verkinto de tiu ĉi Angla poemo estas N. N. 43.

Arg-723-1456 (2012-12-13 22:05:43)

This Irish Christmas Blessing can be found in http: // www. irishcultureandcustoms. com/ Blessings/ Bless2. html .

http://www.poezio.net/version?poem-id=723&version-id=,1456